OpenStudy (yueyue):

Please help me out with this question. cotx sec^4x= cotx + 2tanx + tan ^3x. I have to start from the left side though.

OpenStudy (anonymous):

This requires a lot of steps. I can post a computer solution if you wish.

OpenStudy (yueyue):

Yes please, if you don't mind.

OpenStudy (yueyue):

Whoah, I'm not sure if I understand the solution itself but I'll try to figure it out. Thank you.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

This should be more clear:

OpenStudy (yueyue):

The thing is that I can't really touch the right side of the equation. The question is asking me to verify the right side by substituting identities on the left side.

OpenStudy (anonymous):

I know how to do it. But I don't feel like writing out all the steps myself.

OpenStudy (yueyue):

Ahah that's alright.

OpenStudy (anonymous):

And it's a lot simpler than the two solutions that i've posted.

OpenStudy (campbell_st):

why not start this way cot(x) = cos(x)/sin(x) and sec^4(x) = 1/cos^4(x) so you have cos(x)/sin(x) * 1/cos^4(x) = 1/(sin(x)cos^3(x)

OpenStudy (yueyue):

Yes, alright. Got it. Thank you!